# Optimal map-making with singularities

### 11 Jan 2024 v2

@@ A.2 How to safely reduce the linear system’s size
where $\bm{U}$ has the same shape as $\bm{A}$ and $\bm{U}^{t}\bm{U}$ is an $n_{\rm{pix}}\times n_{\rm{pix}}$ identity matrix, $\bm{D}$ is an $n_{\rm{pix}}\times n_{\rm{pix}}$ diagonal matrix containing the singular values of $\bm{A}$, and $\bm{V}$ is $n_{\rm{pix}}\times n_{\rm{pix}}$ and unitary. Substitute the decomposition into eq. (A.43) shows
@@
where $\bm{U}_{1}$ is $n_{\rm tod}\times n_{\rm tod}$ and unitary; $\bm{D}_{1}$ has $n_{\rm tod}$ rows and $n_{\rm{pix}}$ columns, but most of its elements are zero, except for the $n_{\rm{pix}}\times n_{\rm{pix}}$ square block on top of it, which is diagonal and contains the singular values of $\bm{A}$; and $\bm{V}_{1}$ is still $n_{\rm{pix}}\times n_{\rm{pix}}$ and unitary. Compared with eq. (A.44), this form append normalized and linearly independent columns to the right of $\bm{U}$ to make it unitary, and zeros to the bottom of $\bm{D}$ to match the rule of matrix multiplication, but do not change the matrix multiplication. Thus, the two forms of singular value decomposition are equivalent.
@@
The above equation contains much more rows than eq. (A.45); however, it is easy to observe the following facts:
@@
    The rest lines in eq. (A.47) contain only noise and no information from the sky map.
@@
Items 12 ensure that the sky map information should be solved from the first $n_{\rm{pix}}$ equations, which is exactly eq. (A.45). Item 3 ensures that the noise terms in the rest $n_{\rm tod}-n_{\rm{pix}}$ equations are uncorrelated with the noise terms in the first $n_{\rm{pix}}$ equations; thus, they cannot be used to cancel any noise component in the first $n_{\rm pix}$ rows, i.e., they cannot help to improve the signal-to-noise ratio (SNR). Therefore, the optimal solution given by eq. (A.45, A.47) are identical.
@@ A.3 How to deal with the main matrix’s singularity
Because $\bm{D}_{1}$ is square, diagonal and invertible. It is easy to prove that the optimal solution of the above equation is simply
@@
where $\bm{V}_{1}^{t}$ is a matrix consist of the first $n_{\rm{pix}}-k$ rows of $\bm{V}^{t}$, $\bm{U}_{1}^{t}$ is a matrix consist of the first $n_{\rm{pix}}-k$ rows of $\bm{U}^{t}$, and $k$ is the number of zero rows in eq. (A.50); so we have
@@
In fact, eq. (A.81A.92) are mainly used to derive the solution. Once we obtain eq. (A.92), there is no problem to proceed to the final form of the solution that is free from zero rows or columns. Because both $\bm{U}$ and $\bm{V}$ can be divide in columns as
@@
The above equation give the optimal solution of map-making, provided that the covariance matrix singularity has already been solved by the general approach in section A.1.
@@ A.4 Features of the optimal solution
Provided we have no prior knowledge of the missing modes of the input sky map [^2] The term “no prior knowledge” means the missing modes are regarded as lost forever, and there is no additional knowledge or information that can help to recover them., the optimal solution is the unique solution that satisfy the three important features listed below:
@@ A.5.1 With tightly associated singularities
where $\bm{M}$ is singular. One can prove that in this case, the lower parts of eq. (A.20) are all zero and can be safely ignored to leaves only the upper half
@@
whose noise covariance matrix is non-singular.
In the above mentioned special case, one can prove that the optimal solution is similar to eq.(2.2), just replace each matrix inversion with the pseudo inverse:
@@ A.7 A quick reference
Although the solution in eq. (A.122) may be slightly sub-optimal, it is recommended because it involves a considerably easier computational procedure than the optimal option. Additionally, the divergence from the optimal solution is strictly limited to the singular components, which us usually insignificant.
@@ Appendix B Efficient computation of the noise covariance matrix’s modified pseudo inverse
As long as the filtering scheme does not change, $\bm{M}$ is a constant matrix, so do $\bm{M}^{\times}$ and $\bm{P}^{H}\bm{M}^{\times}$. Therefore, with eqs. (B.6B.7), the time domain eigen-problem is converted to matrix multiplication, which can be computed efficiently once the noise spectrum $\bm{\lambda}$ is obtained.
@@ Appendix C More discussion about the singularity of the main matrix
Thus $\bm{MA}=\bm{U\lambda V}^{t}\bm{P\delta Q}^{t}$. Let $\bm{G}$ be represented as $\bm{V}^{t}\bm{P}$. It is evident that $\bm{G}$ is an $n_{\rm tod}$-by-$n_{\rm tod}$ unitary matrix. Divide $\bm{\lambda}$, $\bm{G}$ and $\bm{\delta}$ into block as
@@
then the only singular part in $\bm{MA}$ can be denote as

### 28 Feb 2024 v3

@@ A.2 How to safely reduce the size of the linear system
where $\bm{U}$ has the same shape as $\bm{A}$ and $\bm{U}^{t}\bm{U}$ is an $n_{\rm{pix}}\times n_{\rm{pix}}$ identity matrix, $\bm{D}$ is an $n_{\rm{pix}}\times n_{\rm{pix}}$ diagonal matrix containing the singular values of $\bm{A}$, and $\bm{V}$ is $n_{\rm{pix}}\times n_{\rm{pix}}$ and unitary. Substituting the decomposition into eq. (A.43) shows
@@
where $\bm{U}_{1}$ is $n_{\rm tod}\times n_{\rm tod}$ and unitary; $\bm{D}_{1}$ has $n_{\rm tod}$ rows and $n_{\rm{pix}}$ columns, but most of its elements are zero, except for the $n_{\rm{pix}}\times n_{\rm{pix}}$ square block on top of it, which is diagonal and contains the singular values of $\bm{A}$; and $\bm{V}_{1}$ is still $n_{\rm{pix}}\times n_{\rm{pix}}$ and unitary. Compared with eq. (A.44), this form appends normalized and linearly independent columns to the right of $\bm{U}$ to make it unitary, and zeros to the bottom of $\bm{D}$ to match the rule of matrix multiplication, but does not change the matrix multiplication. Thus, the two forms of singular value decomposition are equivalent.
@@
The above equation contains many more rows than eq. (A.45); however, it is easy to observe the following facts:
@@
    The remaining lines in eq. (A.47) contain only noise and no information from the sky map.
@@
Items 12 ensure that the sky map information should be solved from the first $n_{\rm{pix}}$ equations, which is exactly eq. (A.45). Item 3 ensures that the noise terms in the remaining $n_{\rm tod}-n_{\rm{pix}}$ equations are uncorrelated with the noise terms in the first $n_{\rm{pix}}$ equations; thus, they cannot be used to cancel any noise component in the first $n_{\rm pix}$ rows, i.e., they cannot help to improve the signal-to-noise ratio (SNR). Therefore, the optimal solution given by eq. (A.45, A.47) are identical.
@@ A.3 How to deal with the main matrix’s singularity
Because $\bm{D}_{1}$ is square, diagonal and invertible, it is easy to prove that the optimal solution of the above equation is simply
@@
where $\bm{V}_{1}^{t}$ is a matrix consisting of the first $n_{\rm{pix}}-k$ rows of $\bm{V}^{t}$, $\bm{U}_{1}^{t}$ is a matrix consisting of the first $n_{\rm{pix}}-k$ rows of $\bm{U}^{t}$, and $k$ is the number of zero rows in eq. (A.50); so we have
@@
In fact, eq. (A.81A.92) are mainly used to derive the solution. Once we obtain eq. (A.92), there is no problem to proceed to the final form of the solution that is free from zero rows or columns. Because both $\bm{U}$ and $\bm{V}$ can be divided in columns as
@@
The above equation gives the optimal solution of map-making, provided that the covariance matrix singularity has already been solved by the general approach in section A.1.
@@ A.4 Features of the optimal solution
Provided we have no prior knowledge of the missing modes of the input sky map [^2] The term “no prior knowledge” means the missing modes are regarded as lost forever, and there is no additional knowledge or information that can help to recover them., the optimal solution is the unique solution that satisfies the three important features listed below:
@@ A.5.1 With tightly associated singularities
where $\bm{M}$ is singular. One can prove that in this case, the lower parts of eq. (A.20) are all zero and can be safely ignored to leave only the upper half
@@
whose the noise covariance matrix is non-singular.
In the above mentioned special case, one can prove that the optimal solution is similar to eq.(2.2), just replacing each matrix inversion with the pseudo inverse:
@@ A.7 A quick reference
Although the solution in eq. (A.122) may be slightly sub-optimal, it is recommended because it involves a considerably easier computational procedure than the optimal option. Additionally, the divergence from the optimal solution is strictly limited to the singular components, which is usually insignificant.
  1. 1.
  2. 2.
  3. 3.
  4. 4.
    When $\bm{A}^{t}\bm{A}$ and $\bm{C}$ are both singular: Can be one of above three.
However, we need to point out the following fact: as mentioned in section 3.1, when the mapmaking algorithm is designed to keep the signal lossless (as a precondition), there is going to be a significant amplification of the noise term in the modes with poor SNR, and the overall mapmaking result may look bad. This is an inevitable consequence of all “lossless” or “make lossless at all costs” mapmaking algorithms. The solution is also simple: one needs to consider a posterior Wiener filtering style approach to put a reasonable balance between the signal loss and noise control.
@@ Appendix B Efficient computation of the noise covariance matrix’s modified pseudo inverse
As long as the filtering scheme does not change, $\bm{M}$ is a constant matrix, so are $\bm{M}^{\times}$ and $\bm{P}^{H}\bm{M}^{\times}$. Therefore, with eqs. (B.6B.7), the time domain eigen-problem is converted to matrix multiplication, which can be computed efficiently once the noise spectrum $\bm{\lambda}$ is obtained.
@@ Appendix C More discussion about the singularity of the main matrix
Thus $\bm{MA}=\bm{U\lambda V}^{t}\bm{P\delta Q}^{t}$. Let $\bm{G}$ be represented as $\bm{V}^{t}\bm{P}$. It is evident that $\bm{G}$ is an $n_{\rm tod}$-by-$n_{\rm tod}$ unitary matrix. Divide $\bm{\lambda}$, $\bm{G}$ and $\bm{\delta}$ into blocks as
@@
then the only singular part in $\bm{MA}$ can be denoted as